OpenStudy (anonymous):

This one is probably a bit interesting: Suppose \(K\) be the number of integers n such that \( \frac{2^n+1}{n^2}\) is also an integer.Find \( K\).

OpenStudy (turingtest):

2^(?)

OpenStudy (anonymous):

write a python script!

OpenStudy (anonymous):

\( \large \frac{2^n+1}{n^2} \)

OpenStudy (anonymous):

for i in [1..infinity] if i 2^n+1/n**2 is an integer: k +=1

OpenStudy (anonymous):

for n*

OpenStudy (anonymous):

n must be odd

OpenStudy (anonymous):

since 2^n + 1 is odd

OpenStudy (anonymous):

no kidding guys, python isn't an option.

OpenStudy (anonymous):

\[2^n + 1 = (2+1)(2^{n-1}-2^{n-2}...-2+1)\]

OpenStudy (anonymous):

Please note this seems to be a hard problem, I only have the the answer, but no concrete proof (yet) or solution.

OpenStudy (anonymous):

So it is finite?

OpenStudy (asnaseer):

I /knew/ it had to be hard as you didn't start it off as "An easy one..." - which are usually quite hard when you set them :-)

OpenStudy (anonymous):

lol, yes it is finite of-course.

OpenStudy (anonymous):

is it the fibonacci sequence?

OpenStudy (anonymous):

asnaseer, you seem be stalking me :D haha

OpenStudy (turingtest):

@agdg that would be infinite

OpenStudy (asnaseer):

:) - you set interesting and challenging questions - thats why

OpenStudy (asnaseer):

OK - I'm off to work on this one and, as Arnie would say, I'll be back!

OpenStudy (anonymous):

does the answer have anything to do with the fibonacci sequence?

OpenStudy (anonymous):

@asnaseer: I told about those 150 right? well the number just moved to 200 so things will get much interesting in the next few weeks ;)

OpenStudy (asnaseer):

FFM - can you please clarify the question - you say K is the NUMBER of integers n such that... did you mean K represents a set of integers n such that? i.e. is K the count of integers or the set of integers satisfying the condition?

OpenStudy (anonymous):

I think the answer is 2 but have no idea how to prove it, that makes it just a guess

OpenStudy (asnaseer):

(2^2+1)/2^2 = 5/4 which is not an integer

OpenStudy (anonymous):

K is the number of n that satisfies the given condition.

OpenStudy (anonymous):

1and 3 does

OpenStudy (anonymous):

n =1 and n = 3 works

OpenStudy (anonymous):

yes why others won't work ?

OpenStudy (anonymous):

I found 2 !

OpenStudy (anonymous):

the answer is K = 2

OpenStudy (anonymous):

but you must show why all the other numbers smaller or bigger than those 2 don't work

OpenStudy (asnaseer):

we need to prove there is no other solution from negative infinity to positive infinity - correct?

OpenStudy (anonymous):

non-positive n is all non-integer

OpenStudy (anonymous):

or am I mistaken?

OpenStudy (asnaseer):

are integers considered to be just the positive numbers?

OpenStudy (anonymous):

I think I got a lead, let me work on a bit.

OpenStudy (anonymous):

yes positive only, I think.

OpenStudy (anonymous):

try plotting the graph for negative integers

OpenStudy (anonymous):

so far I only got K = 2 :(

OpenStudy (asnaseer):

I'm thinking more of the definition of integers - I was always under the impression they were all negative and positive whole numbers

OpenStudy (anonymous):

it only works for 1 and 3 :-D

OpenStudy (anonymous):

we have to prove that \( 2^n+1 \gt n^2 \forall n >3 \)

OpenStudy (anonymous):

that's plain induction routine I suppose ;)

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

alright problem solved give us the next one!

OpenStudy (asnaseer):

I don't get it, how does that prove that \(2^n+1\) is not divisible by \(n^2\)?

OpenStudy (anonymous):

No this is not what you need to prove. Just saying that it is bigger is not enough. It is a divisiblity problem

OpenStudy (anonymous):

it is trivial that is bigger but how do you show taht 2^12589 +1 /12589^2 is not an integer?

OpenStudy (anonymous):

I am really getting slow :P thanks asnaseer :)

OpenStudy (anonymous):

just show that 12589^2 is even

OpenStudy (anonymous):

:) true

OpenStudy (anonymous):

turns out it is odd :-P what do we do now

OpenStudy (asnaseer):

I think I have a lead...

OpenStudy (anonymous):

oh show that n^2 is not a factor of 2^n + 1

OpenStudy (asnaseer):

in one of your previous problems ( http://openstudy.com/users/foolformath#/updates/4f0305f8e4b01ad20b546de4) I proved that \(2^{2k+1}+1\) is always a multiple of 3. I wonder if that can be used here?

OpenStudy (asnaseer):

since here we know n must be odd we can say \(n=2k+1\)

OpenStudy (anonymous):

but, n can be odd or even

OpenStudy (anonymous):

no n must be odd

OpenStudy (asnaseer):

\(2^n+1\) is always odd isn't it?

OpenStudy (anonymous):

yes.

OpenStudy (asnaseer):

therefore \(n^2\) must be odd

OpenStudy (asnaseer):

therefore n must be odd?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aha, right.

OpenStudy (anonymous):

yes?

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

only odd^anything is odd

OpenStudy (anonymous):

2^n + 1 can be expanded as (2 + 1)(2^{n-1} - 2^{n-2} + 2^{n-3}... -2 +1)

OpenStudy (anonymous):

2^n + 1 is always a multiple of 3

OpenStudy (anonymous):

oh so you reduced the domain of n to the odd numbers

OpenStudy (anonymous):

so reduce it to 2 numbers!

OpenStudy (anonymous):

what is the upper bound of the equation

OpenStudy (asnaseer):

so, so we have:\[\frac{2^n+1}{n^2}=\frac{2^{2k+1}+1}{(2k+1)^2}=\frac{3m}{(2k+1)^2}\]

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

?

OpenStudy (asnaseer):

agdgdgdgwngo - I had proved that \(2^{2k+1}=3m\) in an earlier post by FFM

OpenStudy (anonymous):

so 2^2k+1 is a multiple of 3?

OpenStudy (asnaseer):

yes - you can prove it using induction

OpenStudy (anonymous):

so you're close to proving that K = 2

OpenStudy (anonymous):

It can be expanded as (2 + 1)(2^{n-1} - 2^{n-2} + 2^{n-3}... -2 +1). when n is odd

OpenStudy (anonymous):

There is one liner proof that \( 2^n+1\) is a multiple of 3 when n is odd.

OpenStudy (anonymous):

but that uses some number theory notation

OpenStudy (asnaseer):

ok - we know the original numerator \(2^n+1\) is odd, therefore 3m must be odd, therefore m must be odd, so now we have:\[\frac{2^n+1}{n^2}=\frac{2^{2k+1}+1}{(2k+1)^2}=\frac{3m}{(2k+1)^2}=\frac{3(2j+1)}{(2k+1)^2}\]where \(m=2j+1\)

OpenStudy (asnaseer):

aha

OpenStudy (anonymous):

you got it ?

OpenStudy (asnaseer):

this must imply 2j+1 must be divisible by 2k+1 - right?

OpenStudy (asnaseer):

therefore only solutions are when j=k

OpenStudy (asnaseer):

which leaves:\[\frac{3}{2k+1}\]

OpenStudy (asnaseer):

and that only has a solution for k=0 and k=1

OpenStudy (asnaseer):

bingo!

OpenStudy (anonymous):

but I found k = 2

OpenStudy (anonymous):

nice

OpenStudy (asnaseer):

K=2 - I have little k here :-)

OpenStudy (anonymous):

or do you mean that the 2 numbers are 1 and 3

OpenStudy (anonymous):

oh... symbol overload

OpenStudy (asnaseer):

lol!

OpenStudy (anonymous):

my python solution came up with K = 2 for n = 1 and 3 in half a second..... how long did this take? an hour.

OpenStudy (asnaseer):

that felt really satisfying - time to go eat now - thx FFM for setting such wonderful problems - I feel young again! :-)